he slope of the kineticenergy versus positionvecor gives the rate of change of
 

Grade:12th pass

1 Answers

Sunder Muthukumaran
6 years ago
The slope of KE and POSITION VECTOR graph gives the rate of change of MOMENTUM
Here is how , 
KE = 1/2mv2
Position be X 
Then the slope is  KE/x  =  1/2mv÷ x 
=  mv/t (1/2 is a constant (dimensionless), no need to take it to the consideration )
 Therefore , 
 mv/t = rate of change of momentum , which is force .